You need 210 mL of a 90% alcohol solution. On hand, you have a 25% alcohol mixture and a 95% alcohol mixture. How much of each mixture will you need to obtain the desired solution?

Answer by mananth(16949)

percent ---------------- quantity
Alcohol I 95.00% ---------------- x ml
Alcohol II 25.00% ------ 210 - x ml
Mixture 90.00% ---------------- 210
Total 210 ml
95.00% x + 25.00% ( 210 - x ) = 90.00% * 210
95 x + 25 ( 210 - x ) = 18900
95 x + 5250 - 25 x = 18900
95 x - 25 x = 18900 - -5250
70 x = 13650
/ 70
x = 195 ml 95.00% Alcohol I
15 ml 25.00% Alcohol II

m.ananth@hotmail.ca